OpenStudy (epikrebel):

Find the domain of the rational function: g(x)=x/x^2-9

OpenStudy (epikrebel):

\[g(x)=\frac{x}{x^2-9}\]

OpenStudy (anonymous):

the domain is the x-values, so you look at when x does not exist. You cannot divide by zero, so \[x^2-9\neq0\]. you can factor this:\[(x+3)(x-3)\neq0\] so, \[x \neq3\] and \[x \neq-3\] Thus the domain is all the x-values except \[x=3\] and \[x=-3\] In interval notation, that is \[(-\infty,-3)U(-3,3)U(3,\infty)\]

OpenStudy (epikrebel):

I would have also accepted \[{x|x \neq -3, x \neq 3}\] but thank you.